person321:

A compound was analyzed and was found to contain the following percentages of the elements by mass: barium, 89.56%; oxygen, 10.44%. Determine the empirical formula of the compound.

ZombieLover:

) Assume 100 g of the compound is present. Therefore: barium = 89.56 g oxygen = 10.44 g 2) Convert to moles: barium = 89.56 g / 137.33 g/mol = 0.65215 mol oxygen = 10.44 g / 16.00 g/mol = 0.6525 mol 3) Divide by smallest (looking for smallest whole-number ratio): barium = 0.65215 mol / 0.65215 mol = 1 oxygen = 0.6525 mol / 0.65215 mol = 1

person321:

so its BaO?

ZombieLover:

correct :)

person321:

thank you :)

ZombieLover:

no problem

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieLover ) Assume 100 g of the compound is present. Therefore: barium = 89.56 g oxygen = 10.44 g 2) Convert to moles: barium = 89.56 g / 137.33 g/mol = 0.65215 mol oxygen = 10.44 g / 16.00 g/mol = 0.6525 mol 3) Divide by smallest (looking for smallest whole-number ratio): barium = 0.65215 mol / 0.65215 mol = 1 oxygen = 0.6525 mol / 0.65215 mol = 1 \(\color{#0cbb34}{\text{End of Quote}}\) Its better if u don't copy word for word from another user on a different sites answer.